OpenStudy (anonymous):

Find the point on the parabola y=x^2 at which the tangent line is parallel to the line 2x+y=0 a.) (1,1) b.) (0,0) c.) (-1,1) d. (2,4) e.) (-2,4)

OpenStudy (phi):

y=x^2 do you know how to take the derivative? that will give an equation that tells you the slope of the tangent line at any point x

OpenStudy (phi):

also, you need the slope of 2x+y=0 in y = mx +b format you have y = -2x this tells you m = -2 and b=0 m is the slope

OpenStudy (anonymous):

derivative is 2x right?

OpenStudy (phi):

yes. so you want 2x = -2 solve for x then use that value to find y = x^2

OpenStudy (anonymous):

so if slope of 2x+y=0 is -2; and using the derivative of y=x^2 (which is 2x); then my equation should be 2x = -2 (to find the x?) which results in -1 for the point x. substituting that with the original function y=x^2 then the results for y of the point is 1 so the answers is (-1,1) aka letter c?

OpenStudy (anonymous):

oh! thanks i didn't see the last comment. Thank you!